# On the Readiness of Scientific Data for a Fair and Transparent Use in Machine Learning

### v1

@@ Abstract
To ensure the fairness and trustworthiness of machine learning (ML) systems, recent legislative initiatives and relevant research in the ML community have pointed out the need to document the data used to train ML models. Besides, data-sharing practices in many scientific domains have evolved in recent years for reproducibility purposes. In this sense, the adoption of these practices by academic institutions has encouraged researchers to publish their data and technical documentation in peer-reviewed publications such as data papers. In this study, we analyze how this scientific data documentation meets the needs of the ML community and regulatory bodies for its use in ML technologies. We examine a sample of 4041 data papers of different domains, assessing their completeness and coverage of the requested dimensions, and trends in recent years, putting special emphasis on the most and least documented dimensions. As a result, we propose a set of recommendation guidelines for data creators and scientific data publishers to increase their data’s preparedness for its transparent and fairer use in ML technologies.
@@ 1 Introduction
The growing influence of machine learning (ML) technologies in our society raises scientific and political concerns about the potential harms they may cause. Datasets are critical in these systems, and recent research has highlighted them as one of the root causes of unexpected and harmful consequences in ML applications. Recent studies, for example, observed gender-biased classifiers for computer-aided diagnosis due to imbalances in training data [1, 2], or ML models for pneumonia detection failing to generalize to other hospitals due to specific conditions during the collection of the images used to train them [3]. This situation has prompted the interest of regulatory agencies and the ML community in general in developing data best practices, such as building proper dataset documentation. Public regulatory initiatives, such as the European AI ActaaaAI Act required documentation: Annex IV: https://www.euaiact.com/annex/4 and the AI Right of Billsbbbhttps://www.whitehouse.gov/ostp/ai-bill-of-rights/, as well as relevant scientific works [4, 5], have proposed general guidelines for developing standard dataset documentation. Such proposals identify a number of dimensions, such as the dataset’s provenance or potential social issues, that could influence how the dataset is used and the quality and generalization of the ML models trained with it.
@@
The use of this type of data for the training of ML models is making good data practices even more important. However, research has yet to be conducted on the suitability of these data-sharing practices and the data dimensions required by the ML community. In this sense, the goal of this paper is to study the presence (and evolution) of such dimensions in current data papers published by the scientific community. To do so, we analyzed 4041 open access data papers published in two interdisciplinary data journals: Nature’s Scientific Data and Elsevier’s Data in Brief. We examined the full manuscripts regarding the demanded dimensions which are detailed in Section 2. The analysis was aided by an ML pipeline that included a Large Language Model (LLM) at its core, designed to extract specific dimensions from the manuscripts.
@@ 2 Background
The need for proper documentation for datasets in the ML field is well defined in the well-known publication Datasheets for Datasets[4], a work inspired by datasheets in the electronics industry. In this work (among others), the authors identify data aspects for each phase of the data creation process—design, gathering, and annotations—that could affect how the dataset should be used or the quality of ML models trained. In Table 1, we summarize the main dimensions covered in this section.
@@ 3 Results
This section reports on the results of our analysis. We first provide some information on the representativeness of the sample, describing the size and diversity of the data papers, and second, we report on the presence or absence of the information dimensions that characterize a good dataset for ML training purposes. We also provide a temporal evolution of such data to see whether the situation is improving.
3.1 Size and diversity of the sample
The analyzed sample comprises 4041 data papers extracted from two interdisciplinary scientific journals. In Figure 1, we can see an apparent growing trend in the number of papers published between 2015 and 2023, with 2021 being an exception due to the effect of COVID-19 on general scientific production. Note also that the absolute value for 2003 is smaller as we only collected data until June of 2023. The 16,5% of the sample’s data papers are related to people, and this is the subsample we used to calculate the presence of the Social concerns and the Profile of the collection targets dimensions in the next section as these metrics are only relevant for people-related datasets.
As an indication of the diversity of our dataset, Table 3.1 shows the 10 most popular topics of the uses of the data in the analyzed sample, showing the sample diversity. The topics related to the Conservation of biodiversity, RNA sequencing for gene analysis, and Chemistry and Physics are the most common ones. Other topics, such as Breast Cancer and Magnetic Resonance Imaging (MRI), are very interrelated, where most of the datasets around Breast Cancer are composed of MRI images. On the other hand, topics such as COVID-19 have appeared quickly in 2021 and are decreasing over time, as they are strongly correlated with a temporal event such as the coronavirus pandemic. The full results of the topic analysis can be found in our open repository [13].
@@ 3.1 Size and diversity of the sample
Topic
Related data papers (Sample’s percentage)
Example
Conservation of biodiversity 352 (8,7%)
"To study the biodiversity and distribution of fish species, and assess the conservation status of fish species." [21]
RNA sequencing for gene analysis 297 (7,4%)
"Quality control and data analysis in RNA-seq experiments, including quality analysis of raw sequencing data…" [22]
Chemistry and physics 203 (5,0%)
"Soil chemistry analysis, soil classification, and land use analysis in New Zealand’s major agricultural regions." [23]
Breast Cancer 181 (4,5%)
"MRI image processing procedures and constructing breast models for studying the effects of chemotherapy on breast cancer patients." [24]
Magnetic resonance imaging 142 (3,5%)
"Preprocessing of functional MRI data, including motion correction, slice-timing correction, non-brain removal…" [25]
Covid19 104 (2,6%)
"To identify geographies for surveillance and inform global health security…" [26]
Smart Homes 61 (1,5%)
"Gain a deep understanding of residential energy use, analyze household behavior and opinions on energy choices…" [27]
Precipitation 60 (1,5%)
"Comparing different precipitation datasets with observations from GHCN-D 80 for the continental United States except Alaska…" [28]
Data Urban Transport 56 (1,4%)
"Analyzing the impacts of congestion in functional urban areas, benchmarking functional urban areas according to accessibility…" [29]
Cropland 54 (1,3%)
"Optimizing agricultural production strategies and water resources management, as well as validating crop yield…" [30]
Figure 3: Collection: Diversity of types of collection processes

### v2

# On the Readiness of Scientific Data Papers for a Fair and Transparent Use in Machine Learning

@@ Abstract
To ensure the fairness and trustworthiness of machine learning (ML) systems, recent legislative initiatives and relevant research in the ML community have pointed out the need to document the data used to train ML models. Besides, data-sharing practices in many scientific domains have evolved in recent years for reproducibility purposes. In this sense, academic institutions’ adoption of these practices has encouraged researchers to publish their data and technical documentation in peer-reviewed publications such as data papers. In this study, we analyze how this broader scientific data documentation meets the needs of the ML community and regulatory bodies for its use in ML technologies. We examine a sample of 4041 data papers of different domains, assessing their completeness, coverage of the requested dimensions, and trends in recent years. We focus on the most and least documented dimensions and compare the results with those of an ML-focused venue (NeurIPS D&B track) publishing papers describing datasets. As a result, we propose a set of recommendation guidelines for data creators and scientific data publishers to increase their data’s preparedness for its transparent and fairer use in ML technologies.
@@ 1 Introduction
The growing influence of machine learning (ML) technologies in our society raises scientific and political concerns about the potential harms they may cause. Datasets are critical in these systems, and recent research has highlighted them as one of the root causes of unexpected and harmful consequences in ML applications. Recent studies, for example, observed gender-biased classifiers for computer-aided diagnosis due to imbalances in training data [1, 2], or ML models for pneumonia detection failing to generalize to other hospitals due to specific conditions during the collection of the images used to train them [3]. This situation has prompted the interest of regulatory agencies and the ML community in general in developing data best practices, such as building proper dataset documentation. Public regulatory initiatives, such as the European AI Act and the AI Right of Bills, as well as relevant scientific works [4, 5], have proposed general guidelines for developing standard dataset documentation. Such proposals identify a number of dimensions, such as the dataset’s provenance or potential social issues, that could influence how the dataset is used and the quality and generalization of the ML models trained with it.
@@
The use of scientific data published through data papers for the training of ML models is making good data practices even more important. However, research has yet to be conducted on the suitability of these data-sharing practices of the broader scientific community and the data dimensions required by the ML community. In that sense, this paper aims to study the presence (and evolution) of such dimensions in current data papers published by the scientific community. To do so, we analyzed 4041 open access data papers published in two interdisciplinary data journals: Nature’s Scientific Data and Elsevier’s Data in Brief. We examined the full manuscripts regarding the demanded dimensions, which are detailed in Section 2, and we compared the results with an ML-focused venue (NeurIPS Datasets and Benchmarks track) publishing papers describing datasets. The analysis was aided by an ML pipeline that included a Large Language Model (LLM) at its core, designed to extract specific dimensions from the manuscripts.
Using the extracted data, we aimed to answer the following research question:
  • RQ1: To what extent are data papers documented with the dimension demanded by the recent ML documentation frameworks?
    With this question, we aim to see to what extent the data documentation practices outside the ML field already answer the needs stated by the recent ML documentation frameworks.
  • RQ2: How have documentation practices evolved in the last few years regarding the dimension the ML community demands?
    With this question, we aim to see how the recent application of documentation frameworks in the ML community has influenced broader data publishing venues.
@@ 2 Background
The need for proper documentation for datasets in the ML field is well defined in the well-known publication Datasheets for Datasets[4], a work inspired by datasheets in the electronics industry. In this work (among others [14, 15, 16, 3, 17]), the authors identify data aspects for each phase of the data creation process—design, gathering, and annotations—that could affect how the dataset should be used or the quality of ML models trained. Table 1 summarizes the dimensions covered by recent documentation frameworks, representing a common ground between them. This list of dimensions is based on the previous author’s work [5] and the Responsible AI extension of Croissant [18], a metadata language for describing ML datasets promoted by the MLCommons consortium.
@@ 3 Results
This section reports on the results of our analysis. We first provide some information on the representativeness of the sample, describing the size and diversity of the data papers, and second, we report on the presence or absence of the information dimensions that characterize a good dataset for ML training purposes. In this report, we provide an overall picture of the presence of the dimensions, a temporal evolution of such data to see whether the situation is improving, and a comparison between the two analyzed data journals. Then, we present the results of analyzing the data papers of an ML-focused venue (the NeurIPS track on datasets and benchmarks) to draw a picture of the current documentation practices in the ML community and to compare it with the results obtained.
@@
Topic Related data papers (Sample’s percentage) Example
RNA sequencing for gene analysis 371 (9,1%) "Quality control and data analysis in RNA-seq experiments, including quality analysis of raw sequencing data…" [22]
Medical Imaging 330 (8,1%) "MRI image processing procedures and constructing breast models for studying the effects of chemotherapy on breast cancer patients." [23]
Chemistry and physics 232 (5,7%) "Soil chemistry analysis, soil classification, and land use analysis in New Zealand’s major agricultural regions." [24]
Material properties 202 (4,9%) "Analyze and study solubility data, perform high throughput DFT calculations…" [25]
Climate and hydrological analysis 189 (4,6%) "Preprocessing of functional MRI data, including motion correction, slice-timing correction, non-brain removal…" [26]
Biological conservation 160 (3,9%) To study the biodiversity and distribution of fish species, and assess the conservation status of fish species." [27]
Agricultura and cropland 87 (2,1%) "Optimizing agricultural production strategies and water resources management, as well as validating crop yield…" [28]
Drug development 72 (1,8%) "Analyze drug-indication pairs and to study the relationship between drugs and diseases…"[29]
Geological and Seismic analysis 54 (1,4%) reconstructions of paleoclimate data and to analyze the variables of interest for.. [30]
Human movement recognition 53 (1,3%) analyzing locomotor trials and standing trials in healthy subjects and persons with motor disturbances… [31]
@@
3.1 Size and diversity of the sample
The analyzed sample comprises 4041 data papers extracted from two interdisciplinary scientific journals. In Figure 1, we can see an apparent growing trend in the number of papers published between 2015 and 2023, with 2021 being an exception due to the effect of COVID-19 on general scientific production. Note also that the absolute value for 2023 is smaller as we only collected data until June of 2023. The 16,5% of the sample’s data papers are related to people, and this is the subsample we used to calculate the presence of the Social concerns and the Profile of the collection targets dimensions in the next section as these metrics are only relevant for people-related datasets.
As an indication of the diversity of our dataset, Table 3 shows the 10 most popular topics of the uses of the data in the analyzed sample, showing the sample diversity. The topics related to the Conservation of biodiversity, RNA sequencing for gene analysis, and Chemistry and Physics are the most common ones. Other topics, such as Breast Cancer and Magnetic Resonance Imaging (MRI), are very interrelated, where most of the datasets around Breast Cancer are composed of MRI images. On the other hand, topics such as COVID-19 have appeared quickly in 2021 and are decreasing over time, as they are strongly correlated with a temporal event such as the coronavirus pandemic. The full results of the topic analysis can be found in our open repository [13].